(a) Find the real root of the equation z3 + z + 10 = 0 given that one root is 1 − 2i.
(b) Given that 3 + i is a root of the equation z3 − 3z2 − 8z + 30 = 0, find the
remaining roots.
(c) Given that 1 + i is a root of the equation z3 − 2z + k = 0, find the other two
roots and the value of the real constant k.
(d) Given that 2 − 3i is a root of the equation z3 + pz2 + qz + 13 = 0, find the other
two roots and the values of the real constants p and q.
(e) Show that z = i is a root of the equation z4 + z3 + z − 1 = 0. Find the three
other roots.
1
Expert's answer
2020-05-21T16:26:31-0400

a) let f(z) = z3+z+10=0


If z=1-2i is a root of f , then, z=1+2i is also a root of f as complex roots occur in conjugate pairs.

thus, z = 1-2i and z= 1+2i


z-1 = "\displaystyle \mp" i


(z-1)2+4 divides f(z)


f(z) = [(z-1)2+4] [ z+2]


the real root of f(z) is z= -2

verification: f(-2) = 0


b) let f(z) = z3 − 3z2 − 8z + 30 = 0


If z=3+i is a root of f , then, z=3-i is also a root of f as complex roots occur in conjugate pairs.

thus, z = 3+i and z= 3-i


(z-3)2 + 1 divides f(z)


f(z) = [(z-3)2+1] [ z+3]


the real root of f(z) is z= -3

verification: f(-3) = 0


c) let f(z) =  z3 − 2z + k = 0


If z=1+i is a root of f , then, z=1-i is also a root of f as complex roots occur in conjugate pairs.

thus, z = 1+i and z= 1-i


(z-1)2 + 1 divides f(z), so remainder will be equal to zero


f(z) = [(z-1)2+1] [ z+2] + (k-4)


here, the remainder is (k-4) = 0

thus, k=4


f(z) = z3 − 2z + 4


the real root of f(z) is z= -2

verification: f(-2) = 0



d) let f(z) =  z3 + pz2 + qz + 13 = 0

If z=2-3i is a root of f , then, z=2+3i is also a root of f as complex roots occur in conjugate pairs.

thus, z = 2-3i and z= 2+3i


(z-2)2 + 9 divides f(z), so remainder will be equal to zero


f(z) = [(z-2)2+9] [ z+(4+p)] + [(q-13)z-4(4+p)z+13-13(4+p)]


here, the remainder is [(q-13)z-4(4+p)z+13-13(4+p)]= 0

thus, p=-3 , q=9


f(z) = z3-3z2+9z+13


the real root of f(z) is z= -1

verification: f(-1) = 0


e) let f(z) = z4+z3+z-1=0


If z=i is a root of f , then, z=-i is also a root of f as complex roots occur in conjugate pairs.

thus, z = i and z= -i


(z)2+1 divides f(z)


f(z) = [(z)2+1] [ z2+z-1]


the real roots of f(z) is z= "\frac{-1\displaystyle \mp \sqrt{5}}{2}"


verification: f("\frac{-1+ \sqrt{5}}{2}" ) = 0 and f("\frac{-1- \sqrt{5}}{2}" ) =0